OpenStudy (anonymous):

Evaluate r+s/8 r=44 and s = 20 simplify your answer

OpenStudy (saifoo.khan):

44 + 20/8

OpenStudy (anonymous):

44+20 /8 64/8 =8

OpenStudy (saifoo.khan):

93/2

OpenStudy (anonymous):

does it look like this? \[\frac {r+s}{8}\]

OpenStudy (anonymous):

Saifoo is that simplified?

OpenStudy (anonymous):

jo math yes it does

OpenStudy (saifoo.khan):

Oh Yeah!

OpenStudy (saifoo.khan):

in mixed fractions: 46 1/2

OpenStudy (anonymous):

Then its (44+20)/8 = 64/8 = 8

OpenStudy (anonymous):

The answer is 8.

OpenStudy (anonymous):

Saifoo that is simplified too? Do you think I should write it as 46.5?

OpenStudy (saifoo.khan):

hey, is your ques looks like this? (44+20)/8 or 44 + (20/8)

OpenStudy (anonymous):

(44+20)/8

OpenStudy (saifoo.khan):

Oh, then the answer is "8".

OpenStudy (anonymous):

ok thank you !!!!

OpenStudy (saifoo.khan):

uw. :)